Status Report

NASA Space Station On-Orbit Status 23 January 2006

By SpaceRef Editor
January 23, 2006
Filed under , , ,
NASA Space Station On-Orbit Status 23 January 2006

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2006) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below.  Underway: Week 16 for Expedition 12.

The crew is on shifted sleep cycle schedule this week for accommodating to next weeks EVA-15 spacewalk (which begins at ~5:26pm EST). The shift began on 1/21 by 3 hours (wake up at 4:00am) and added another 2 hours yesterday, to 6:00am EST (sleep time: 9:30pm), where it will remain until after the EVA.

Early in the day, Bill McArthur and Valery Tokarev had 4 hours reserved to prepare for the spacewalk by studying flight procedures, a training DVD and the preliminary EVA timeline. The review was supported by ground specialist tagup and supplemented by more tagups with EVA-15 specialists via S-band/audio.

Tokarev broke out and set up the equipment for tomorrow s planned Russian Urolux biochemical urine test (PZE MO-9).  [MO-9 is conducted regularly every 30 days (and also before and after EVAs) and is one of five nominal Russian medical tests adopted by NASA for US crewmembers for IMG PHS (Integrated Medical Group/Periodic Health Status) evaluation as part of the “PHS/Without Blood Labs” exam. The analysis uses the sophisticated in-vitro diagnostic apparatus Urolux developed originally for the Mir program. The data are then entered in the Medical Equipment Computer (MEC) s special IFEP software (In-Flight Examination Program).]

The FE relocated the ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS) payload hardware from the DC1 Docking Compartment to the Service Module (SM) and activated the experiment.  [The AST spectrometer and its kits, including the ALS return pouch and kits with materials (ESKILO) were stowed in the portside crew cabin (crew cabin 1). The ESA/RSC-Energia experiment ALTCRISS uses the AST spectrometer employed by VC8 guest cosmonaut Roberto Vittori last year in the DC1 for the Italian LAZIO (Low Altitude Zone Ionization Observatory) experiment. Progress 20 brought up a new shielding belt made of Nomex, containing polyethylene bricks and two new dosimeters in a dedicated pocket.]

Using the gear gathered yesterday by Tokarev for water sampling, the CDR today collected KAV condensate water samples upstream of the Gas-Liquid Mixture Filter (FGS) of the SRVK-2M condensate water processor into a temporarily installed KAV humidity condensate container (#6577), for return to Earth.

McArthur transferred the backup SODF (Station Operations Data File) DVD from Progress 20, thus completing another task from his post-20P unpack list, and configured the CD Library (Vol. II) for transfer to the Russian Segment (RS) when the hatches are closed in preparation for the spacewalk from the RS.

Bill also readied the Nikon F5 digital EVA camera with the 28-mm lens for the spacewalk, leaving it connected to vehicle power in the Lab until it is moved to the RS.

Later, the CDR continued preparing the EVA tools required for the external activities.

Tokarev completed the periodic (about monthly) replenishing of the Elektron s water supply for electrolysis, his fifth, filling the KOV thermal loops EDV container with purified (deionized) water from the BKO multifiltration/purification column unit, following carefully written instructions.  [The procedure is specially designed to prevent air bubbles larger than ~20 mm from getting into the new BZh-8 Liquid Unit where they could cause Elektron shutdown. In the procedure, the EDV water is carefully drawn from the BKO and the air/liquid separator unit (GZhS) while the crewmember checks for any air bubbles in the EDV (and, if visible, estimates their number, with up to 10 bubbles of less than 20 mm diameter permitted). Elektron water is also supplied from USOS condensate in a CWC (collapsible water container) that is checked for its contents of air bubbles and is rejected if the estimated total air bubble volume is more than 30 cubic centimeters (1 cm air bubble is about 0.5 ccm).]

Valery performed the daily routine maintenance of the Service Module (SM)’s environment control & life support system (SOZh), while Bill, working off his discretionary task list, completed the weekly collection of the toilet flush (SP) counter and water supply (SVO) readings for calldown to TsUP/Moscow.

Processing Status
Daily Mission
Return to Flight
Weekly Status
Weekly Science
Daily On-Orbit Status
Daily Crew Timeline
Soyuz | Progress
ISS News | ATV

Both crewmembers completed their regular 2.5-hr. physical exercise program on the TVIS treadmill, RED resistive exerciser and VELO bike with bungee cord load trainer.  [Valery s daily protocol prescribes a strict four-day microcycle exercise with 1.5 hr on the treadmill in unmotorized mode and one hour on VELO plus load trainer (today: Day 2 of the first set).]

Afterwards, the CDR transferred the exercise data files to the MEC (Medical Equipment Computer) for downlink, as well as the daily wristband HRM (heart rate monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

At ~7:00pm EST, the crew is scheduled for their weekly 15-min. teleconference with ISS Program Management at JSC/Houston via S-band/audio.

The Science Officer continued his work on the PromISS-4 (Protein Crystal Growth Monitoring by Digital Holographic Microscope #4) experiment, today removing video tape #12 from the MSG (Microgravity Science Glovebox) video system.

The ground-commanded BCAT-3 (Binary Colloidal Alloy Test) activity is continuing, taking time-lapse flash photography of BCAT sample 6 at the MWA (Maintenance Work Area) via EarthKAM camera and SSC-7 laptop. McArthur conducted a check of the alignment and focus of the camera on the sample and position of flash. The imaging is to continue until 1/26.

Working off his discretionary time available task list, Tokarev worked on the BIO-5 Rasteniya-2 (“Plants-2”) experiment, completing the daily status check and the regular periodic download of data & imagery collected of the experiment to the computer for subsequent downlink to the ground.  [Rasteniya researches growth and development of plants under spaceflight conditions in the Lada-8 greenhouse. The regular maintenance of the experiment (each Monday, Wednesday, Friday and Sunday) involves monitoring of seedling growth, humidity measurements, moistening of the substrate if necessary, topping off the water tank if ~20-25% of the total amount (4 liters) remains, and photo/video recording. Once weekly, data from the Lada greenhouse control unit are recorded on floppy disk for weekly downlink via REGUL-Packet or the new BSR-TM at a suitable occasion

Valery also checked the operation of the Japanese experiment GCF-JAXA (Granada Crystallization Facility) in the Russian TBU incubator, maintained at 20 degC, including a temperature check on its ART (automatic temperature recorder).  [This daily monitoring/temp checking, carried on the Russian voluntary “time available” task list, will continue until 4/30.]

The current week will be dominated by more EVA preparations onboard. The Orlan spacewalk from the DC1 airlock on 2/3 will last an estimated 5h 58m, beginning at 5:26pm EST and ending at 11:08pm.  [Its major objectives are: Deployment of RadioSkaf (an old Orlan spacesuit equipped with ham radio equipment for radio amateurs on Earth), removal of the Russian/US Strela 2 from the EFGF (electrical flight grapple fixture) on the FGB module and its installation on the FRGF (flight releasable grapple fixture) on the PMA-3 (pressurized mating adapter #3), retrieval of the Biorisk experiment, and various external inspections and photographing.]

No CEO (Crew Earth Observation) photo targets uplinked today.

Over 177,000 of CEO (Crew Earth Observation) images have been taken in the first five years of the ISS.

  CEO photography can be viewed and studied at the websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 12 crew visit:

Expedition 12 Flight Crew Plans can be found at

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

ISS Orbit  (as of this morning, 6:55am EST [= epoch]):

  • Mean altitude — 347.0 km
  • Apogee height — 353.7 km
  • Perigee height — 340.2 km
  • Period — 91.48 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.0010036
  • Solar Beta Angle — -71.0 deg (magnitude peaking tomorrow)
  • Orbits per 24-hr. day — 15.74
  • Mean altitude loss in last 24 hours — 60 m
  • Revolutions since FGB/Zarya launch (Nov. 98) — 41020

Events Ahead (all dates Eastern; tentative):

  • 02/03/06 — Russian EVA-15
  • 02/11/06 — ISS Reboost Test (in MMOD avoidance mode)
  • 03/03/06 — Progress M-54/19P undocking & reentry
  • 03/30/06 — Soyuz TMA-8/12S launch (Exp. 13 + Marcus Pontes/Brazil)
  • 04/01/06 — Soyuz TMA-8/12S docking (DC1)
  • 04/24/06 — Progress M-56/21P launch
  • 04/26/06 — Progress M-56/21P docking
  • 06/19/06 — Progress M-55/20P undocking & reentry

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.